OpenStudy (anonymous):

Cheyenne drove 3000 miles in 75 hours. At this rate how long will it take her to drive 4000 miles. I need work to go with the answer And work that I can do. I'm only in 6th grade

OpenStudy (anonymous):

3000x=300000 x=300000/3000 x=100 hrs

OpenStudy (aripotta):

set up a proportion\[\frac{ 3000~mi }{ 75~hrs }=\frac{ 4000~mi }{ x~hrs }\]

OpenStudy (aripotta):

cross-multiply and solve for x

OpenStudy (nicholausblackmon):

AriPotta!

OpenStudy (nicholausblackmon):

Here to whip her wand and complete math problems in a second!

OpenStudy (aripotta):

lol :)